# Testing identifying assumptions in Tobit Models

### v1

@@ 3.2. IV Tobit model
We turn our attention to the Instrumental Variable Tobit case and propose testable implications that can be used to test the model described in 2.10 and the associated assumptions 3 and 4.
The observed data includes $(Y,D,Z)$, and we characterise the joint distribution of $(Y,D)\in\mathcal{Y}\times\mathcal{D}$ conditional on the instrument, $Z$, to obtain the model’s testable implications. The mapping from observed probabilities to its corresponding quantities implied by the IV Tobit model requires jointly evaluating the continuous (interior) support for the outcome and treatment variables, as well as the accumulation points and distribution tails conditional on $Z$. Consider any $0<c_{0}<c_{1},d_{0}<d_{1}$, and $z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$,
@@
$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z;\rho)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.7)
@@
$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (3.9)
@@
where $\Phi_{W,V}(w,v;\rho)$ denotes the joint c.d.f. of $(W,V)$, a bivariate normal with coefficient of correlation $\rho$. Similarly, considering the case that the observed $Y$ equals zero,
@@
In the context of model 2.10 or 2.16, equalities 3.4-3.12 for all $c,d,z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$ are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},\tilde{D},\tilde{U},\tilde{V},Z)$ or equivalently $(\tilde{Y},\tilde{D},\tilde{W},\tilde{V},Z)$ that satisfies model 2.10 and 2.16, Assumptions 3 and 4, and induces the observed distribution on the data $(Y,D,Z)$. See Appendix B for the proof.
Then, the equalities 3.4-3.12 are sharp testable implications for the validity of the Instrumental Variable Tobit Model in 2.10 coupled with assumptions 3-4. They serve as the basis for the test procedure described in Section 5. A similar claim of non-learnability holds for the iv-tobit. The demonstration follows a similar logic.
@@ 4.1. Classic Tobit model
The coefficients have been normalized by dividing by the latent variable’s ($U$) square root of the variance, denoted by $\sigma$. Naturally, this implies that the ratio of the original parameters, say $\tilde{\alpha}_{0},\tilde{\alpha}_{1}$ and $\sigma$ is identified since $P(Y=0|D=d)=\Phi(\alpha_{0}+\alpha_{1}D)=\Phi(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D)$.
Similarly to the case analysed by Han and Vytlacil (2017), this allows the identification of the inverse mills ratio which is a known function of the standard normal density, $\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$. Then,
@@ 5. Testing procedure
To test the sharp equalities, we rewrite each of these equalities as two inequalities (Mourifié and Wan, 2017; Acerenza, Bartalotti, and Kédagni, 2023). For a concrete example, we first note that for example
$P(c_{1}\leq Y|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\iff E[1\{c_{1}\leq D\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]=0.$
We then we rewrite each of these moment equalities implied by the restrictions on the empirical distribution as moment inequalities
$E[1\{c_{1}\leq D\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0,E[-1\{c_{1}\leq D\}+1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0.$
In doing this, for all the restrictions on the empirical distribution we can implement a test relying on existing intersection bounds inferential methods such as Chernozhukov, Lee, and Rosen (2013) which is specifically suited to test conditional moment inequalities.
Note that the equalities for the classic Tobit need to hold for any pair of constants $(c_{0},c_{1})$, and the ones from the IV Tobit need to hold for pairs of $(c_{0},c_{1})$ and $(d_{0},d_{1})$. We propose a partition of $\mathcal{Y}\times\mathcal{D}$ to test sufficient conditions of these sharp equalities. Rejection of any of the null hypotheses that the equalities hold at these particular levels will imply violations of the model.
@@
where $\hat{\theta}(D,k)$ is a nonparametric estimator for $\theta(D,k)$, $\hat{s}(D,k)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
@@ IV Tobit model
For the instrumental variable Tobit model, the continuous support for both the outcome and treatment poses challenges to the implementation of the test. The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ and $d_{0},d_{1}\in\mathcal{D}$ are given by equations 3.4-3.12.
@@
$W_{kq}=\left\{\begin{array}[]{cc}1\{Y=0\}1\{D\leq d_{0}\}-\Phi^{2}_{d_{0}},&\text{ if }k=0,q=0\\ 1\{Y=0\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{2}_{d_{q},d_{q+1}},&\text{ if }k=0,1\leq q<Q\\ 1\{Y=0\}1\{D\geq d_{Q}\}|-\Phi^{2}_{d_{Q}},&\text{ if }k=0,q=Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{k+1},c_{k},d_{0}},&\text{ if }1\leq k<K,q=0\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{k+1},c_{k},d_{q+1},d_{q}},&\text{ if }1\leq k<K,1\leq q<Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{k+1},c_{k},d_{Q}},&\text{ if }1\leq k<K,q=Q\\ 1\{Y\geq c_{K}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{K},d_{0}},&\text{ if }k=K,q=0\\ 1\{Y\geq c_{K}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{K},d_{q+1},d_{q}},&\text{ if }k=K,1\leq q<Q\\ 1\{Y\geq c_{K}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{K},d_{Q}},&\text{ if }k=K,q=Q.\\ \end{array}\right.$
@@
where $\hat{\theta}(D,k)$ is a nonparametric estimator for $\theta(D,k)$, $\hat{s}(D,k)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
@@
Intuitively, the testable conditions derived above consider if the empirical conditional distribution of the observed outcome variable - in both the mass accumulation and non-truncated parts of the support of $Y$ is consistent with a random variable(s) following the (bivariate) normal distribution for different sections of the distribution and values of the independent instrument, $Z$.
@@
3. Violations due to the empirical distributions of the residuals differing from the implied distributions in certain values of the treatment (Classic Tobit) or instrument (IV Tobit), which indicate violations of the exogeneity of treatment or instrument (kedagni_generalized_2020).
Remark 6 (On the inclusion of covariates).
@@
The existing procedure can be extended to include exogenous covariates, $X$, within the linear index model in 2.10. Then, the testable equalities for both classic and IV Tobit models could be derived with the additional conditioning on $X$. To implement the test with covariates, we suggest generating a partition of the covariate space, say in $J$ grids, similarly to the partition of the exogenous variable used in Section 5. For every grid in the partition, one could compute the test statistic 5.3 or 5.6 for the classic or the IV-tobit, respectively. Then obtain critical values that account for multiple testing via a Bonferroni correction. In particular, for every grid, set the critical value at the significance level $\frac{\alpha}{J}$, namely $\kappa_{1-\frac{\alpha}{J}}$.
A theoretically interesting approach would partition the joint support of all the exogenous variables and compute the test statistics across all parts of the grid. However, this approach would entail significant implementation challenges, since the statistical packages for Chernozhukov et al. (2015) allow for only one exogenous covariate.
@@ 6. Simulations
In this section, we provide simulation exercises for the size and power of the proposed test for the classical Tobit model assumptions.
The testing procedure described relies on testing sharp equalities that should hold for any arbitrary partition. Since both the outcome and treatment are continuous, evaluating all possible equalities is technically challenging. We focus on a non-sharp set of the equalities by evaluating them at different grid partitions of their support, in a similar spirit to honore2020selection. Naturally, this choice makes the test less powerful as we don’t consider the continuum of equalities derived in Section 4, but is justified by the ease of implementation of the procedure based on intersection bounds and the performance of the test on the simulations below.
For the simulations presented below, we considered a partition of the support of the observed outcome variable into the accumulation point ($Y=0$) and four quartiles on the (untruncated) positive range. The choice of the number and location of the partitions/evaluation points balances the implementation computational requirements, availability of data for different parts of the joint support of the outcome, treatment and exogenous instrumental variable. For the procedure in Section 5 to be feasible, we must have data on both the outcome and the exogenous variables within each partition. Using the empirical quantiles of the non-truncated outcomes to determine the partitions guarantees a reasonable number of observations for each part of the grid. Larger sample sizes might allow finer partitions for the outcome support to be used, but the added computational requirements created by an increased number of equalities being checked, coupled with the larger datasets, can substantially increase computing time. [^7] The simulations presented in Section 6 have limited sample sizes and a relatively small number of equalities being tested due to the long running time and computational constraints when repeating the test procedure thousands of times.
@@
6.1. Classic Tobit model
6.1.1. Size
For the classic Tobit model, we consider a continuous treatment $D$ following the data generating process described in Equation 6.8.
@@ 6.1.1. Size
Where $\mathbf{0}_{p}$ is a $p\times 1$ vector of zeroes. The parameter $\rho$ determines the intensity of dependence between the latent variables jointly determining the treatment and outcome. Under the treatment exogeneity condition described in Assumption 1, $\rho=0$. Table 1 presents the empirical test sizes for this scenario, under different significance levels $\alpha$. The results indicate that while the test over rejects the null hypothesis for small to mid-sized samples, the test’s empirical coverage approaches its desired nominal benchmark as samples larger than 5,000 are used.
Table 1. Test Size
@@
6.1.2. Power
To consider the test’s performance under violations of the exogeneity assumption, we use the DGP described in Equation 6.8 with different values for $\rho$, reflecting various degrees of endogeneity. Larger values of $\rho$ produce more acute violations of the null hypothesis. Table 2 presents results for these scenarios: columns 3-4 present the MSE while the remaining columns present the empirical rejection rates. As expected, the power of the test increases with larger $\rho$ and bigger sample sizes, presenting substantial power for $\alpha=10\%$ and presenting a decreasing pattern for $\alpha=5\%$ and $\alpha=1\%$.
Table 2. Test power for violations in exogeneity

### v2

@@ 3.2. IV Tobit model
We turn our attention to the instrumental variable Tobit case and propose testable implications that can be used to test the model described in 2.8 and the associated assumptions 3 and 4.
The observed data includes $(Y,D,Z)$, and we characterize the joint distribution of $(Y,D)\in\mathcal{Y}\times\mathcal{D}$ conditional on the instrument, $Z$, to obtain the model’s testable implications. The mapping from observed probabilities to their corresponding quantities implied by the IV Tobit model requires jointly evaluating the continuous (interior) support for the outcome and treatment variables, as well as the accumulation points and distribution tails conditional on $Z$. Consider any $0<c_{0}<c_{1},d_{0}<d_{1}$, and $z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$,
@@
$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$ (3.7)
@@
$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)$ (3.9)
@@
where $\Phi_{W,V}(w,v;\rho)$ denotes the joint c.d.f. of $(W,V)$, a standard bivariate normal with coefficient of correlation $\rho$. Similarly, considering the case that the observed $Y$ equals zero,
@@
In the context of model 2.8 or 2.12, equalities 3.4-3.12 for all $c,d,z\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}$ are sharp in the sense that whenever they hold, it is possible to construct a vector of $(\tilde{Y},\tilde{D},\tilde{U},\tilde{V},Z)$ or equivalently $(\tilde{Y},\tilde{D},\tilde{W},\tilde{V},Z)$ that satisfies model 2.8 and 2.12, Assumptions 3 and 4, and induces the observed distribution on the data $(Y,D,Z)$. See Appendix B for the proof.
Remark 5 (Non-learnability).
The testable implications and sharpness discussed above show that the IV-Tobit model is generally refutable. However, the model is non-verifiable. The demonstration follows a similar logic as in the classic Tobit case. See Appendix E for the proof.
Then, the equalities 3.4-3.12 are sharp testable implications for the validity of the instrumental variable Tobit Model in 2.8 coupled with assumptions 3-4. They serve as the basis for the test procedure described in Section 5.
@@ 4.1. Classic Tobit model
The coefficients have been normalized by dividing by the latent variable’s ($U$) square root of the variance, denoted by $\sigma$. Naturally, this implies that the ratio of the original parameters is identified since $P(Y=0|D=d)=\Phi(\alpha_{0}+\alpha_{1}D)=\Phi(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D)$. This allows the identification of the inverse mills ratio, which is a known function of the standard normal density, $\lambda(\tilde{\alpha}_{0}/\sigma+\tilde{\alpha}_{1}/\sigma D))$. Then,
@@ 5. Testing procedure
To test the sharp equalities, we rewrite each of these equalities as two inequalities (Mourifié and Wan, 2017; Acerenza, Bartalotti, and Kédagni, 2023). For a concrete example, we first note that
$P(c_{1}\leq Y|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)\iff E[1\{c_{1}\leq Y\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]=0.$
We rewrite each of these moment equalities implied by the restrictions on the empirical distribution as moment inequalities
$E[1\{c_{1}\leq Y\}-1+\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0,E[-1\{c_{1}\leq Y\}+1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}D)|D]\leq 0.$
In doing this for all the restrictions on the empirical distribution, we can implement a test relying on existing intersection bounds inferential methods such as Chernozhukov, Lee, and Rosen (2013), which is specifically suited to test conditional moment inequalities.
Note that the equalities for the classic Tobit hold for any pair of constants $(c_{0},c_{1})$, and the ones from the IV Tobit hold for pairs of $(c_{0},c_{1})$ and $(d_{0},d_{1})$. We propose a partition of $\mathcal{Y}\times\mathcal{D}$ to test sufficient conditions of these sharp equalities. Rejection of any of the null hypotheses that the equalities hold at these particular levels implies violations of the model.
@@
where $\hat{\theta}(D,k)$ is a nonparametric estimator for $\theta_{k}(D)$, $\hat{s}(D,k)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
@@ IV Tobit model
For the instrumental variable Tobit model, the continuous support for both the outcome and treatment poses challenges to the implementation of the test. [^8] Note that a discrete treatment can also be accommodated as an intermediate step between the current derivation and the derivations from Acerenza, Bartalotti, and Kédagni (2023). The sharp testable equalities for every $c_{0},c_{1}\in\mathcal{Y}$ and $d_{0},d_{1}\in\mathcal{D}$ are given by equations 3.4-3.12.
@@
$W_{kq}=\left\{\begin{array}[]{cc}1\{Y=0\}1\{D\leq d_{0}\}-\Phi^{2}_{d_{0}},&\text{ if }k=0,q=0\\ 1\{Y=0\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{2}_{d_{q},d_{q+1}},&\text{ if }k=0,1\leq q<Q\\ 1\{Y=0\}1\{D\geq d_{Q}\}-\Phi^{2}_{d_{Q}},&\text{ if }k=0,q=Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{k+1},c_{k},d_{0}},&\text{ if }1\leq k<K,q=0\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{k+1},c_{k},d_{q+1},d_{q}},&\text{ if }1\leq k<K,1\leq q<Q\\ 1\{c_{k}\leq Y\leq c_{k+1}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{k+1},c_{k},d_{Q}},&\text{ if }1\leq k<K,q=Q\\ 1\{Y\geq c_{K}\}1\{D\leq d_{0}\}-\Phi^{1}_{c_{K},d_{0}},&\text{ if }k=K,q=0\\ 1\{Y\geq c_{K}\}1\{d_{q}\leq D\leq d_{q+1}\}-\Phi^{1}_{c_{K},d_{q+1},d_{q}},&\text{ if }k=K,1\leq q<Q\\ 1\{Y\geq c_{K}\}1\{D\geq d_{Q}\}-\Phi^{1}_{c_{K},d_{Q}},&\text{ if }k=K,q=Q.\\ \end{array}\right.$
@@
where $\hat{\theta}(Z,k,q)$ is a nonparametric estimator for $\theta_{kq}(Z)$, $\hat{s}(D,k,q)$ its standard error, and $\kappa_{1-\alpha}$ is a critical value at the significance level $\alpha$.
@@
Intuitively, the testable conditions derived above consider whether the empirical conditional distribution of the observed outcome variable — in both the mass accumulation and non-truncated parts of the support of $Y$ — is consistent with a random variable(s) following the (bivariate) normal distribution for different sections of the distribution and values of the independent instrument, $Z$.
@@
3. Violations due to the empirical distributions of the residuals differing from the implied distributions in certain values of the treatment (Classic Tobit) or instrument (IV Tobit), which indicate violations of the exogeneity of treatment or instrument (Kédagni and Mourifié, 2020).
Including Covariates
@@
The proposed procedure can be extended to include exogenous covariates, $X$, within the linear index model in 2.8. Then, the testable equalities for classic and IV Tobit models can be derived with the additional conditioning on $X$. The test with covariates could be implemented by generating a partition of the covariate space, say in $J$ grids, similar to the partition of the exogenous variable used in Section 5. For every grid in the partition, one computes the test statistic 5.3 or 5.6 for the classic or the IV-tobit, respectively. Then obtain critical values that account for multiple testing via a Bonferroni correction. In particular, for every grid, set the critical value at the significance level $\frac{\alpha}{J}$, namely $\kappa_{1-\frac{\alpha}{J}}$.  [^9] A theoretically interesting approach would partition the joint support of all the exogenous variables and compute the test statistics across all parts of the grid. However, this approach would entail significant implementation challenges since the statistical package for Chernozhukov et al. (2015) allows for only one exogenous covariate. This procedure can be cumbersome when there are many covariates or when they are continuous.
Another route that is less computationally intensive follows Acerenza, Bartalotti, and Kédagni (2023). We illustrate it for the $IV$-tobit case, and a similar logic could be applied to the classic Tobit case. Thus, with covariates added to the linear index function, we have,
$\left\{\begin{array}[]{lcl}Y&=&\max(0,Y^{*})\\ Y^{*}&=&\beta_{0}+\beta_{1}Z+\beta_{2}X+W\\ D&=&\gamma_{0}+\gamma_{1}Z+\gamma_{2}X+V\end{array}\right.$ (5.10)
We extend Assumption 3 to formalize covariates’ exogeneity:
Assumption 5 (full independence).
[…]
Under Assumptions 4-5, by a similar argument to that used in Section 3, any given generic testable implication becomes
$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z,X=x)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
for all $z\in\mathcal{Z}$ and $x\in\mathcal{X}$, where $x$ can be a vector. Equivalent conditions for the case that $Y=0$ and other parts of the support of $Y$ and $D$ can be similarly obtained.
Implementation
Since the test is carried out nonparametrically, we face challenges when the dimension of $X$ is large, especially with continuous covariates. Another operational limitation is that the Stata’s clrbound package only allows for one conditioning variable at a time. For those practical reasons, we propose implementing a weaker, non-sharp, version of the testable equalities. Continuing with the example of Equality 5, we can integrate over the covariates $X$ (or alternatively, $Z$), taking advantage of the fact that $\mathbb{E}[W|Z=z,X=x]=0$ implies $\mathbb{E}[W|Z=z]=0$ and $\mathbb{E}[W|X=x]=0$ for all random variables $W$.
Implementation becomes very similar to the IV-Tobit discussed above by redefining the simplifying notation in equations (3.4)-(3.12), with the only difference being the inclusion of $X$ on the linear indexes in $\Phi_{W}(\cdot)$ and $\Phi_{W,V}(\cdot,\cdot,\cdot)$. For example,
$\Phi^{1}_{c_{1},c_{0},d_{1},d_{0}}$$\equiv$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{1}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z-\beta_{2}x,d_{0}-\gamma_{0}-\gamma_{1}z-\gamma_{2}x;\rho),$
and similarly for the other terms. Then, compute the new $W_{kq}$ in the same manner as in Section 5, and the intersection bounds framework considers the similar $2(K+1)(Q+1)$ inequalities, based on the partition of the support of $Y$ and $D$.
$\sup_{z}E[W_{kq}|Z=z]$$\leq$ 00 0
$\sup_{z}E[-W_{kq}|Z=z]$$\leq$ $0,\text{ for }k=0,\dots,K;q=0,\dots,Q.$
This is the test procedure that we implement for the empirical application in Section 8.
Under Assumption 5, one could base the test in conditioning on a particular covariate $X_{C}$ instead of $Z$ by integrating the sharp Equality 5 over $Z$ and then implementing the intersection bounds procedure using $X_{C}$ as the sole conditioning variable. [^10] Assumption 5 constraints the relationship between the covariates and the latent error terms in such a way that one could combine the information on all $X$ and $Z$ in an index and construct non-sharp testable equalities that would hold conditional on this index. For example, having $\sup_{t}E[W_{kq}|Z+X_{1}+...+X_{C}=t]=0$, for $k=0,\dots,K;q=0,\dots,Q$.
@@ 6. Simulations
In this section, we provide simulation exercises for the proposed tests. The testing procedure described relies on testing sharp equalities that should hold for any arbitrary partition of the outcome, treatment and instrument support. When both the outcome and treatment are continuous, evaluating all possible equalities is technically challenging. We focus on a non-sharp set of the equalities by evaluating them at different grid partitions of their support, in a similar spirit to Honoré and Hu (2020) which is particularly well suited for continuous supports. Naturally, this choice makes the test less powerful as we don’t consider the continuum of equalities derived in Section 4, but is justified by the ease of implementation of the procedure based on intersection bounds and the performance of the test on the simulations below.
For the simulations related to the classic Tobit, we partition the support of the observed outcome variable into the accumulation point ($Y=0$) and four quartiles on the (untruncated) positive range, while for the IV-Tobit case we also partition the support of the treatment variable ($D$) to create a grid based on both the outcome and treatment. [^11] Note that if the treatment variable was discrete, the respective grid points could be set naturally to the possible countable values $D$ takes. The choice of the number and location of the partitions/evaluation points balances the implementation computational requirements, data availability for different parts of the joint support of the outcome, treatment and exogenous instrumental variable. For the procedure in Section 5 to be feasible, we must have data on both the outcome and the exogenous variables within each partition. Using the empirical quantiles of the non-truncated outcomes to determine the partitions guarantees a reasonable number of observations for each grid part. Larger sample sizes might allow finer partitions for the outcome support. Still, the added computational requirements created by an increased number of equalities being checked, coupled with the larger datasets, can substantially increase computing time. [^12] The simulations presented in Section 6 have limited sample sizes and a relatively small number of equalities being tested due to the long-running time and computational constraints when repeating the test procedure thousands of times.
@@
In evaluating the finite sample performance of the proposed tests, we consider a continuous treatment $D$ following the data generating process described in Equation 6.8.
@@
Where $\mathbf{0}_{p}$ is a $p\times 1$ vector of zeroes. The parameter $\rho$ determines the intensity of dependence between the latent variables jointly determining the treatment and outcome. Under the treatment exogeneity condition described in Assumption 1, $\rho=0$. Table 1 presents the empirical test sizes for the Classic Tobit test this scenario, for different significance levels $\alpha$. The results indicate that while the test over rejects the null hypothesis for small to mid-sized samples, the test’s empirical coverage approaches its desired nominal benchmark as samples larger than 5,000 are used.
Table 1. Classic Tobit Test Size
@@
To consider the test’s performance under violations of the exogeneity assumption, we modify the DGP in Equation 6.8 with different values for $\rho$, reflecting various degrees of treatment endogeneity. Larger values of $\rho$ produce more acute violations of the null hypothesis. Table 2 presents the results for the Classic Tobit test: columns 3-4 report the MSE, while the remaining columns show empirical rejection rates. As expected, the power of the test increases with larger $\rho$ and bigger sample sizes. [^14] Simulations for test performance when violations of normality of the latent errors occur are presented in Appendix F.
Table 2. Classic Tobit test power for violations in exogeneity